OpenStudy (anonymous):

PLZ HELP ME!!! You work as a cashier for a grocery store and earn $5 per hour. You also mow lawns and earn $10 per hour. You want to earn at least $50 per week, but would like to work no more than 10 hours per week. Which system of inequalities, along with y greater than or equal to 0 and x greater than or equal to 0, would you use to solve the real-world problem?

OpenStudy (anonymous):

you work at grocery story (x) and you work mowing lawns (y)....you want to work no more then 10 hrs... so your first inequality will deal with hrs. x + y <= 10 (that a less then or equal to sign...you do not want to work more then 10 hrs, but you can still work 10 hrs, so we include the equal sign as well) grocery store (x) earning $5 per hr and lawn mowing (y) earning $10 per hr...making at least $50... now we will make an inequality dealing with the $ 5x + 10y >= 50 (that a greater then or equal to sign...you want to make 50 or more, so we include the equal sign) so basically, your 2 inequalities are : x + y <= 10 (you do not have to use g and m...you can use any variable you wish) 5x + 10y >= 50